OpenStudy (anonymous):

19 missylyn1990 0Suppose x coins are tossed. Write an expression to represent the number of possible outcomes.

OpenStudy (anonymous):

This looks like a online assignment question -.-. it has a problem number ......

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

its a copy/paste from an online assignment, but there is no way to connect to an online school; so its not deletable. Just bewary on how an answer is presented :)

OpenStudy (anonymous):

smh. either way there are 2 possible outcomes. tada :D

OpenStudy (anonymous):

now how do you represent those?

OpenStudy (anonymous):

its from ashworth community.com from a test im tryin to do

OpenStudy (anonymous):

Suppose x coins are tossed. Write an expression to represent the number of possible outcomes. A. 2x B. x+x C. x.x D. 2x

OpenStudy (anonymous):

@amistre64 even test is said. this is obvious cheating

OpenStudy (anonymous):

im not really takeing the test m trying to understand how to do it

OpenStudy (anonymous):

so can u plz help me i really dont understand any of this

OpenStudy (amistre64):

number of possible outcomes reads vague to me. It sounds like a probability question but it reads terribly to me

OpenStudy (amistre64):

spose we have 4 coins hhhh hhht hhth hthh thhh hhtt 1 4 6 4 1 is the pattern that will emerge for all possible outcomes; giving us a total of 16 if 3 coins are tossed: 1 3 3 1 = 8 outcomes; so x*x aint it and x+x aint it. 2*x aint it either; so if one of those is an exponent, that might be the working solution. 2^4 = 16 2^3 = 8 etc ....

OpenStudy (amistre64):

1 5 10 10 5 1 = 32 = 2^5 so yeah, i think my idea might have merit

OpenStudy (anonymous):

thank u i understand know (: